OpenStudy (anonymous):

Use the quadratic formula to find the zeros of the function. Round to tenths if necessary. y = 17x2 - 21 A. {-1.1, 1.1} B. {1.1} C. {-4.1, 4.6} D. no real solution Use the quadratic formula to find the zeros of the function. Round to tenths if necessary. y = 17x2 - 21 A. {-1.1, 1.1} B. {1.1} C. {-4.1, 4.6} D. no real solution @Mathematics

OpenStudy (anonymous):

\[17x^2-2=0\] \[x^2=\frac{2}{17}\] \[x=\pm\sqrt{\frac{2}{17}}\] then a calculator